# Semantic Consistent Language Gaussian Splatting for Point-Level Open-vocabulary Querying

### v1

@@ 4.1 Constructing consistent language supervision
14:     ${\mathcal{B}}\leftarrow\texttt{GetBoxFromMask}({\mathbb{S}}_{{\bm{I}}_{t}})$
15:     $\tilde{\mathbb{S}}_{1:T}\leftarrow\tilde{\mathbb{S}}_{1:T}\cup\texttt{SAM2}({\bm{I}}_{1:T},{\mathcal{B}})$
@@
Constructing consistent ground-truth. With the set of masklets $\tilde{\mathbb{S}}_{1:T}$ extracted, we create an average CLIP embedding $\bar{\phi}_{r}$ for each masklet $\tilde{\mathcal{S}}_{r}\in\tilde{\mathbb{S}}_{1:T}$. This is done by masking out the image ${\bm{I}}_{t}$ using the extracted masklet and then passing it to CLIP’s image encoder, i.e.,
@@
Intuitively, as averaging reduces variance, the proposed $\bar{\phi}_{r}$ is more consistent than the individual $\phi_{r,t}$ used in LangSplat. Furthermore, the weighting scheme helps to reduce the influence of small regions that often contain noisier language embeddings. In contrast, a naive equal weight to each observation from different views disregards the reliability of individual features.
@@
GSGroup. [44] 7.75 8.80 10.94 16.29 10.95 10.71 9.86 10.17 27.27 14.50 10.71 2.82 5.08 4.55 5.79
Ours 58.98 37.85 43.57 29.67 42.52 82.14 61.97 54.24 50.00 62.09 30.36 9.86 11.86 13.64 16.43
Table 1: Quantitative results on the LERF dataset. For OpenGaussian, we report the numbers from their paper.
@@
Table 2: Quantitative results on the 3D-OVS dataset.
Figure 4: Qualitative results on LERF dataset of scene“ramen” and “figurines”. For each scene, the first row contains rendered language embeddings, and the second row contains 3D query results.
@@
Table 3: Quantitative results on Replica dataset. Note, to compute mIoU and mAcc using the ground-truth point clouds, we skip the densification stage when training 3D Gaussian Splatting for all methods.
@@
As shown, for pixels across views that are tracked and segmented into the same region, this construction of ground-truth assigns the same averaged CLIP embedding. Compared with LangSplat’s ground-truth in Eq. (5), our construction is shared across time $t$, i.e., the language embedding ${\bm{l}}_{i}$ in LangSplat receives consistent supervision across all relevant frames.
Figure 5: Qualitative results on 3D-OVS dataset for scene “lawn”. The first row contains rendered language embeddings, and the second row contains 3D query results for “green lawn”.
Autoencoder details. As in LangSplat, to reduce the GPU memory usage, we train a light-weight autoencoder consisting of an encoder E and a decoder D. Differently, we encode the averaged CLIP embedding $\bar{\phi}_{r}$ into a low-dimensional latent space, i.e.,$\texttt{E}(\bar{\phi}_{r})$ is used as supervision to train the language embeddings ${\bm{l}}_{i}$.
4.2 Two-step retrieval for 3D Querying
With the text query vector ${\bm{q}}$, the standard one-step approach directly compares the CLIP features ${\bm{q}}$ of the query text with the language embeddings ${\bm{l}}_{i}$ of each Gaussian, i.e.
^$\hat{\mathcal{M}}_{\bm{q}}^{\tt 3D,1step}\triangleq\{{\bm{g}}_{i}|~{}\forall i~{}\texttt{Cos}({\bm{l}}_{i},{\bm{q}})\geq\text{threshold}\},$ (13)
where Cos is the cosine similarity. However, this one-step approach struggles to find a single effective threshold across different language embedding ${\bm{l}}_{i}$ as CLIP’s image and language embedding are known to be not well calibrated [20]. To address this challenge, we propose our two-step approach for querying 3D Gaussians, specifically designed to mitigate this problem.
Step-1. Given the CLIP feature of a text query ${\bm{q}}\in{\mathbb{R}}^{512}$, we first retrieve the most similar average feature over all regions feature
$\bar{\phi}^{*}_{r}\triangleq\operatorname*{arg\,max}_{\bar{\phi}_{r}}\texttt{Cos}(\bar{\phi}_{r},{\bm{q}}).$ (14)
Since $\bar{\phi}_{r}$ is obtained as a weighted average of CLIP image embeddings and ${\bm{q}}$ comes from CLIP text embeddings, a direct comparison between them through cosine similarity is effective and does not involve a threshold.
Step-2. With the retrieved average feature $\bar{\phi}^{*}_{r}$, we leverage the pretrained encoder E to compress it into lower dimension. Then we compute its cosine similarity with the learned language embedding ${\bm{l}}_{i}$ for each Gaussian. Next, we threshold this similarity to retrieve the tentative set of Gaussians
~$\tilde{\mathcal{M}}_{\bm{q}}^{\tt 3D}=\{{\bm{g}}_{i}|~{}\forall i~{}\text{Cos}({\bm{l}}_{i},\texttt{E}(\bar{\phi}^{*}_{r}))\geq\text{threshold}\},$ (15)
where the querying process compares $\texttt{E}(\bar{\phi}^{*}_{r})$ with ${\bm{l}}_{i}$. Recall $\texttt{E}(\bar{\phi}^{*}_{r})$ is used as supervision for training language embeddings ${\bm{l}}_{i}$, i.e., the relevant language embeddings are trained to be similar to $\texttt{E}(\bar{\phi}^{*}_{r})$. Therefore, any high threshold works well, which improves the queries’ reliability and robustness. More details are provided in the Appendix Sec. A4.
Finally, we apply DBSCAN [8] as post-processing to this retrieved set to remove outliers based on the spatial position $\mu_{i}$ of the Gaussians.
5 Experiments
Datasets. Following LangSplat [28], we conduct experiments on the further annotated LERF [16] dataset that contains a set of in-the-wild scenes and on the 3D-OVS [23] dataset, which includes a collection of long-tail objects for evaluating open-vocabulary 3D querying. Additionally, we report results on the Replica [37] dataset, which has labeled point clouds for indoor objects. We randomly choose 10 classes as text queries from the 8 chosen scenes of Replica.
@@ 5 Experiments
Baselines. We consider three baselines:
  • LangSplat [28], which also trains a language 3D Gaussian. Hence, the 1-step query approach in Eq. (13) can be directly used, following how OpenGaussian [40] evaluates.
  • OpenGaussian [40] proposed the task of open-vocabulary 3D (point-level) querying. As they reported on the LERF dataset, we directly included their results from the paper. For 3D-OVS and Replica datasets, we use their publicly released implementation.
  • GaussianGrouping [44] We follow their implementation for the open-vocabulary querying based on the first frame’s semantics.
@@
Implementation details. To extract language features, we use the OpenCLIP ViT-B/16 model. For 2D mask segmentation, we employ the SAM ViT-H model, and for tracking masks of the same object, we utilize the SAM2-hiera-large model. Other hyperparameters regarding 3D Gaussians follow LangSplat. More details are in the Appendix Sec. A3.
5.1 Quantitative results
@@
LERF dataset. In Tab. 1, we show the results on the LERF dataset. We observe that Ours consistently outperforms LangSplat and, on average, is better than OpenGaussian, achieving an improvement of $+4.16$ in mIoU and $+10.66$ in mAcc. We observe that LangSplat and GaussianGrouping, methods designed for 2D querying, do not generalize well to point-level querying.
@@
3D-OVS dataset. The results for the 3D-OVS dataset is reported in Tab. 2. Our method achieves 70.58% in mIoU and 93.33% in Loc. Acc, significantly outperforming baseline methods. Notably, our method demonstrates a $+20.42$ gain in mIoU averaged across scenes. We observe that there exists $100\%$ in the Loc. Acc because the dataset is relatively easy with $\leq 30$ frames and $\leq 7$ objects in each scene, leading to effective segmentation and tracking from SAM2. Additionally, we observed that OpenGaussian has a low Loc. Acc due to small “floaters” in the retrieved Gaussians. These floaters shift the center point of the exterior bounding box to an incorrect location.
@@
Replica dataset. We show experimental results in Tab. 3. We observe that our approach outperforms OpenGaussian with an average gain of $+1.74$ on mIou and $+9.29$ on mAcc. There exists $0.00$ in GaussianGrouping’s results because their implementation only uses the first frame’s semantics. For large scenes such as Replica, the query object may not appear in the first frame, leading to an empty query.
5.2 Qualitative results
@@
3D-OVS dataset. Fig. 5 shows the qualitative results on 3D-OVS dataset. Our method creates more consistent language embeddings with crisp object boundaries. Furthermore, our method’s retrieval results are less noisy and preserve the complete structure of each queried object, i.e., the lawn is accurately retrieved with query “green lawn”.
Replica dataset. Fig. 6 visualizes the queried points for a scene in the Replica dataset given the query “cloth”. We observe that LangSplat and GaussianGrouping failed to retrieve the correct object, and OpenGaussian only retrieves part of the cloth with noisy points from other objects. Overall, our method retrieved a more complete and clearer object, which is consistent with the quantitative result.
5.3 Ablation study.
We conduct ablation studies to validate the efficacy of each proposed component of our method and report the performance in Tab. 4. Recall, $\bar{\phi}_{r}$ denotes the weighted average features. Without SAM2’s masklets to get consistent features of the region, the performance drops significantly. Also, the 2-step query helps to get more accurate retrieval in 3D space. Overall, we observe that all proposed components are necessary to achieve the optimal performance.
We also conduct ablation studies on the effectiveness of our method without DBSCAN and evaluate the performance of canonical querying from LERF [16] in 3D querying. The original definition of canonical query from LERF is conducted on rendered 2D images. To evaluate the performance of canonical query in 3D querying, we replace $\phi_{\texttt{lang}}$ by ${\bm{l}}_{i}$ in the original equation. Formally, for each language embedding ${\bm{l}}_{i}$ and each text query ${\bm{q}}$, the relevancy score is defined as $\min_{i}\frac{\exp({\bm{l}}_{i}\cdot{\bm{q}})}{\exp({\bm{l}}_{i}\cdot{\bm{q}})+\exp({\bm{q}}\cdot\phi_{canon}^{i})}$ where $\phi_{canon}^{i}$ is the CLIP embedding of a predefined canonical phrase chosen from “object”, “things”, “stuff”, and “texture”.
Next, we show ablations on canonical query in Tab. 5. We see that without DBSCAN, our proposed method also gains improvement in the performance. We also see that our proposed 2-step query significantly outperforms the canonical query.
Ablations Performance Metrics
SAM2 $\bar{\phi}_{r}$ 2-Step mIoU mAcc Loc. Acc
@@
48.84 67.86 23.21
9.09 12.50 8.93
58.98 82.14 30.36
Table 4: Ablation study on each proposed component using LERF’s “figurines” scene. Note, the average feature $\bar{\phi}_{r}$ cannot be implemented without SAM2. The best result is achieved with all our proposed components.
@@
SAM2 2-Step DBSCAN Cano. Query mIoU mAcc Loc. Acc
@@
Table 5: Ablation study on DBSCAN and Canonical Query.
6 Conclusion
@@
We study the task of open-vocabulary 3D understanding formulated as a point-level 3D querying task. Based on LangSplat’s framework, we propose to improve the ground truth supervision to train more consistent language parameters by leveraging masklets extracted from SAM2. Furthermore, we introduce a novel 2-step querying pipeline to address the difficulty of choosing a consistent threshold in the baseline’s single-step query. Experiments over three datasets demonstrate that our approach achieves state-of-the-art performance.
Appendix
The appendix is organized as follows:
  • In Sec. A1, we discuss the rationale behind the experiment setup, particularly why we did not report on the ScanNet dataset.
  • In Sec. A2, we provide more qualitative results on different datasets.
  • In Sec. A3, we provide our implementation details.
  • In Sec. A4, we provide more analysis on the necessity of the two-step query.
Figure A1: The original 3D Gaussian Splatting produces low-quality 3D reconstruction on the ScanNet dataset released from OpenGaussian.
Query “dark cup” “wavy noodles” “tea in a glass”
RGB Refer to caption Refer to caption Refer to caption
LangSplat Refer to caption Refer to caption Refer to caption
OpenGaussian Refer to caption Refer to caption Refer to caption
GaussianGrouping Refer to caption Refer to caption Refer to caption
Ours Refer to caption Refer to caption Refer to caption
Figure A2: More qualitative results of text query for LERF dataset.
Query “pikachu” “stapler” “banana”
RGB Refer to caption Refer to caption Refer to caption
LangSplat Refer to caption Refer to caption Refer to caption
OpenGaussian Refer to caption Refer to caption Refer to caption
GaussianGrouping Refer to caption Refer to caption Refer to caption
Ours Refer to caption Refer to caption Refer to caption
Figure A3: More qualitative results of text query for 3DOVS dataset.
Query “cloth” “sofa” “blanket”
Ground Truth Refer to caption Refer to caption Refer to caption
LangSplat Refer to caption Refer to caption Refer to caption
OpenGaussian Refer to caption Refer to caption Refer to caption
GaussianGrouping Refer to caption Refer to caption Refer to caption
Ours Refer to caption Refer to caption Refer to caption
Figure A4: More qualitative results of text query for Replica dataset. Yellow points represent the queried objects.
A1 Experimental Design Rationale
For a fair comparison with the concurrent work OpenGaussian, we wanted to have a comparison using the ScanNet dataset as done in their paper. However, we found that their released dataset contains broken camera transformation matrices. Specifically, in the transforms_train.json file of scene0200_00 and scene0400_00. Critically, using the provided data, we observed low-quality 3D reconstructions when using 3D Gaussian Splatting, as illustrated in Fig. A1. Therefore, we did not report on ScanNet as the 3D reconstruction setup is already subpar, and the comparison would be focused on the language feature component. Hence, for a third dataset, we chose the Replica dataset to run our experiments which resulted in more accurate 3D reconstruction for titall methods. In this case, the comparison is more meaningful and focused on the language and query components.
A2 Additional Qualitative Results
We provide additional figures showing the qualitative results. Also attached with this appendix are HTML files showing videos containing additional qualitative comparisons.
More query results. Fig. A2, Fig. A3 and Fig. A4 show more object query visualizations from different datasets. As can be seen, our queried objects are more complete and precise across various queries.
Query results shown in video. We provide HTML with videos visualizing the queried objects’ views from different angles. Please see the file query_comparison/query_comparison.html These visualizations further demonstrate our retrieved objects are more precise and complete than those retrieved by the baseline methods.
3D point features visualization. Besides the rendered features shown in the main paper, we now show,  Fig. A6, the 3D point features of our method on LERF and 3DOVS datasets. Generally, the features are consistent on the same object, which is beneficial for retrieving objects using our 2-step query pipeline.
Video visualization of consistent semantic ground truth. To showcase the ground-truth supervision constructed from our approach, we provide videos illustrating the ground-truth feature ${\bm{L}}_{t}$ used to supervise the training of language embeddings. Please see gt_consistency/gt_consistency.html. We observe that LangSplat’s semantic features often struggle with inconsistency when the same object is viewed from different angles. In contrast, our approach produces more consistent and stable semantic features. This consistency benefits training by reducing ambiguity in supervising the same object across varying viewpoints, resulting in more robust and reliable learned language embeddings.
A3 Implementation Details
(a) Foundation model specification. To extract language features, we use the OpenCLIP ViT-B/16 model. For 2D mask segmentation, we employ the SAM ViT-H model, and for tracking masks of the same object, we utilize the SAM2-hiera-large model. During evaluation, our trained model is assessed with SAM level 2 on the LERF dataset and SAM level 3 on the 3D-OVS and Replica datasets.
(b) Training strategy. For the lightweight autoencoder, we adopt a learning rate of 0.0006 for SAM level 2 and 0.008 for SAM level 3. The structure of the encoder is [256, 128, 128, 64, 32, 3], and decoder is[16, 32, 64, 128, 256, 256, 512]. For 3D scene reconstruction, we follow LangSplat’s approach by pretraining the standard 3D Gaussian Splatting for 30,000 steps. This is followed by training the language embeddings for an additional 30,000 steps, skipping the densification stage. Experiments are conducted on NVIDIA A100 GPU.
(c) Replica dataset. We select the following scenes from the Replica dataset: office_0, office_1, office_2, office_3, office_4, room_0, room_1, room_2. To be consistent with OpenGaussian’s evaluation metrics, we skip the densification stage when training the original 3D Gaussian Splatting on the Replica dataset. The 10 categories defined by Replica and used for text queries are: ceiling, sofa, cloth, stool, blanket, blinds, cushion, pot, rug, picture.
(e) Query thresholding. During querying, we retain only the points with a cosine similarity to the queried text features exceeding a specified threshold. The threshold is set to 0.995 for LERF, 0.999 for 3DOVS, and 0.95 for the Replica dataset.
Figure A5: IoU metric per query vs. cosine similarity thresholds. We observe that the 1-step method does not have a suitable threshold for all queries, while our 2-step method does not face this issue.
A4 More Analysis on 2-step Query Necessity
In Fig. A5, we visualize the IoU Metric for each of the queries in the scene bench in 3D-OVS. We observe that the threshold required for optimal performance varies across different queries. For example, the pebbled wall needs a threshold of 0.4, whereas the dressing doll needs a threshold of 0.6. On average, there is no single threshold that maximizes the performance across all queries. In contrast, our proposed 2-step query does not suffer from this problem. All queries benefit from using a threshold close to 1.
RGB Reference 3D Feature Visualization
Refer to caption Refer to caption
Refer to caption Refer to caption
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure A6: Feature visualization our approach in 3D.

### v2

@@ IV-A Constructing consistent language supervision
14:   ${\mathcal{B}}\leftarrow\texttt{RegionFromMask}({\mathbb{S}}_{{\bm{I}}_{t}})$
15:   $\tilde{\mathbb{S}}_{1:T}\leftarrow\tilde{\mathbb{S}}_{1:T}\cup\texttt{Track}({\bm{I}}_{1:T},{\mathcal{B}})$
@@
Extracting masklets for consistent semantics. Reviewed in Sec. III, a tracking module takes a sequence of images and regions of interest as input to track masks of the same region. In LangSplat, SAM is used to extract the initial candidates, i.e., for each frame ${\bm{I}}_{t}$, SAM proposes a set of regions ${\mathbb{S}}_{{\bm{I}}_{t}}$. Starting from the first frame, we check if a proposed region has already been tracked by comparing the mIoU of the SAM mask with the tracked masks and applying a threshold. If the proposed region has not been tracked, we run the tracking model and add the output masklets to the set of tracked masklets $\tilde{\mathbb{S}}_{1:T}$. These steps are summarized in Alg. 1.
Constructing consistent ground-truth. With the set of masklets $\tilde{\mathbb{S}}_{1:T}$ extracted, we create a consistent ground-truth by aggregating the CLIP embedding $\bar{\phi}_{r}$ for each masklet $\tilde{\mathcal{S}}_{r}\in\tilde{\mathbb{S}}_{1:T}$. This is done by masking out the image ${\bm{I}}_{t}$ using the extracted masklet and then passing it to CLIP’s image encoder:
@@
The main intuition behind this design is that averaging reduces variance. The proposed $\bar{\phi}_{r}$ is more consistent than the individual $\phi_{r,t}$ used in LangSplat when used as supervision for the distillation. Furthermore, the weighting scheme helps to suppress the contribution of small regions that often contain noisier language embeddings, i.e., we consider the reliability of individual features.
TABLE I: Quantitative results on LERF. For OpenGaussian, we report the numbers from their paper.
@@
Ours 58.91 37.85 43.57 29.67 42.50 82.14 61.97 54.24 50.00 62.09 82.14 61.97 62.71 40.91 61.93
TABLE II: Quantitative results on the 3D-OVS dataset.
@@
TABLE III: Quantitative results on Replica dataset. To compute mIoU and mAcc using the ground-truth point clouds, we skip the densification stage when training 3D Gaussian Splatting for all methods.
@@
For pixels across views that are tracked and segmented into the same region, this construction of ground-truth assigns the same averaged CLIP embedding. Compared with LangSplat’s ground-truth in Eq. (5), our construction is shared across time $t$, i.e., the language embedding ${\bm{l}}_{i}$ in LangSplat receives consistent supervision across all relevant frames.
Figure 5: Qualitative results on LERF dataset of scene“ramen" and “figurines". For each scene, the first row contains rendered language embeddings, and the second row contains 3D query results.
Figure 6: Qualitative results on 3D-OVS dataset for scene “lawn". The first row contains rendered language embeddings, and the second row contains 3D query results for “green lawn".
@@
IV-B Ground Truth (GT)-anchored 3D Querying
With the text query vector $\bm{q}$}, the standard approach is to directly compares the CLIP features $\bm{q}$} of the query text with the language embeddings ${\bm{l}}_{i}$ of each Gaussian, i.e.
$\hat{\mathcal{M}}_{\bm{q}}^{\tt 3D,1step}\triangleq\{{\bm{g}}_{i}|~\forall i~\texttt{Cos}({\bm{l}}_{i},{\bm{q}})\geq\text{threshold}\},$ (10)
where Cos is the cosine similarity. However, this one-step approach struggles to find a single effective threshold across different language embeddings ${\bm{l}}_{i}$ as CLIP’s image and language embeddings are known to be not well calibrated [15]. To address this challenge, we propose our GT-anchored approach for querying 3D Gaussians. This involves first retrieving the ground truth and then comparing the similarity relative to the ground truth, as described in more detail below.
GT retrieval. Given the CLIP feature of a text query ${\bm{q}}\in{\mathbb{R}}^{512}$, we first apply a low threshold to filter out invalid prompts. We then retrieve the most similar average feature (GT for distillation) over all regions feature
$\bar{\phi}^{*}_{r}\triangleq\operatorname*{arg\,max}_{r\in\{r^{\prime}|\texttt{Cos}(\bar{\phi}_{r}^{\prime},{\bm{q}})\geq\text{threshold}\}}\texttt{Cos}(\bar{\phi}_{r},{\bm{q}}).$ (11)
As $\bar{\phi}_{r}$ is obtained as a weighted average of CLIP image embeddings and $\bm{q}$} comes from CLIP text embeddings, a direct comparison between them through cosine similarity is effective and does not involve a threshold.
Relative comparison to GT. With the retrieved GT $\bar{\phi}^{*}_{r}$, we compress it into lower dimension with the pretrained encoder E. Then we compute its cosine similarity with the learned language embedding ${\bm{l}}_{i}$ for each Gaussian. Next, we threshold this similarity to retrieve the tentative set of Gaussians
$\tilde{\mathcal{M}}_{\bm{q}}^{\tt 3D}=\{{\bm{g}}_{i}|~\forall i~\text{Cos}({\bm{l}}_{i},\texttt{E}(\bar{\phi}^{*}_{r}))\geq\text{threshold}\},$ (12)
where the querying process compares $\texttt{E}(\bar{\phi}^{*}_{r})$ with ${\bm{l}}_{i}$. Recall $\texttt{E}(\bar{\phi}^{*}_{r})$ is used as supervision for training language embeddings ${\bm{l}}_{i}$, i.e., the relevant language embeddings are trained to be similar to $\texttt{E}(\bar{\phi}^{*}_{r})$. Therefore, any high threshold works well, which improves the queries’ reliability and robustness.
V Experiments
Datasets. Following LangSplat [23], we conduct experiments on the further annotated LERF [12] dataset that contains a set of in-the-wild scenes and on the 3D-OVS [18] dataset, which includes a collection of long-tail objects for evaluating open-vocabulary 3D querying. Additionally, we report results on the Replica [28] dataset, which has labeled point clouds for indoor objects. We then evaluate on ten object classes for querying over eight senses in Replica.
Evaluation metrics. As the ground-truths for LERF and 3D-OVS datasets are provided in 2D, we measure the performance of 3D querying indirectly. We render the queried Gaussians in $\hat{\mathcal{M}}_{\bm{q}}^{\tt 3D}$ to obtain a 2D mask for evaluation using 2D metrics following LangSplat [23]. This includes mean Intersection over Union (mIoU $\uparrow$) and localization accuracy (Loc. Acc $\uparrow$). Here, mIoU is defined as the ratio of the intersecting pixels to the total number of pixels in the union of the predicted masks and ground-truth masks.
[…]
Baselines. For a fair comparison, we strictly followed OpenGaussian [29] for the task of open-vocabulary 3D (point-level) querying. As they reported on the LERF dataset, we directly included their results from the paper. For 3D-OVS and Replica datasets, we use their publicly released implementation. To further benchmark the performance, we included more baseline methods modified (m) for direct 3D querying: LangSplat-m [23], which also trains a language 3D Gaussian. Hence, the standard query approach in Eq. (10) can be directly used, following how OpenGaussian [29] evaluates. GaussianGrouping-m [34], which we follow their implementation for the open-vocabulary query to select group IDs, and use the corresponding Gaussian points as candidates.
@@ V Experiments
Implementation Details. To extract language features, we use the OpenCLIP ViT-B/16 model. For 2D mask segmentation, we employ the SAM ViT-H model, and for tracking masks of the same object, we utilize the SAM2-hiera-large model. Pretraining the standard 3D Gaussian Splatting takes 30,000 steps. This is followed by training the language embeddings for an additional 30,000 steps, skipping the densification stage. Experiments are conducted on an NVIDIA A100 GPU.
V-A Quantitative results
@@
LERF dataset. In Tab. I, we show the results on the LERF dataset. We observe that Ours consistently outperforms LangSplat-m and, on average, is better than OpenGaussian, achieving an improvement of +4.14+4.14 in mIoU and +10.66+10.66 in mAcc. We believe this gain is significant, as retraining the language Gaussians 5 times with different seeds has a small standard deviation, e.g., on figurines the std is 0.24. We observe that LangSplat-m and GaussianGrouping-m, methods designed for 2D querying, do not generalize well to point-level querying.
@@
3D-OVS dataset. The results for the 3D-OVS dataset is reported in Tab. II. Our method achieves 70.58% in mIoU and 93.33% in Loc. Acc, significantly outperforming baseline methods. Notably, our method demonstrates a +20.42+20.42 gain in mIoU averaged across scenes. We observe that there exists $100\%$ in the Loc. Acc because the dataset is relatively easy with $\leq 30$ frames and $\leq 7$ objects in each scene, leading to effective segmentation and tracking from SAM2. Next, we observed that OpenGaussian has a low Loc. Acc due to small “floaters” in the retrieved Gaussians. These floaters shift the center point of the exterior bounding box to an incorrect location.
@@
Replica dataset. We show results in Tab. III. We observe that our approach outperforms OpenGaussian with an average gain of +1.74+1.74 on mIoU and +9.29+9.29 on mAcc. There exists $0.00$ in GaussianGrouping-m’s results because their query implementation only uses the first frame’s semantics. In large scenes like Replica, the query object may not appear initially, causing an empty query.
V-B Qualitative results
@@
3D-OVS dataset. Fig. 6 shows the qualitative results on 3D-OVS dataset. Our method creates more consistent language embeddings with crisp object boundaries. Furthermore, our method’s retrieval results are less noisy and preserve the complete structure of each queried object, i.e., the lawn is accurately retrieved with query “green lawn".
Replica dataset. Fig. 7 visualizes the queried points for a scene in the Replica dataset given the query “cloth”. We observe that LangSplat-m and GaussianGrouping-m failed to retrieve the correct object, and OpenGaussian only retrieves part of the cloth with noisy points from other objects. Overall, our method retrieved a more complete and clearer object, which is consistent with the quantitative result.
TABLE IV: Ablation study on each proposed component using LERF’s “figurines” scene. Note that combining the language features relies on leveraging tracking information and is not feasible without it. So we put "-" in the first row. The best result is achieved with all our proposed components.
Ablations Metrics
Tracking $\bar{\phi}$ Strategy GT-anchored mIoU mAcc Loc. Acc
@@
$\bar{\phi}_{r}$ 9.09 12.50 8.93
$\bar{\phi}_{r}^{\texttt{stand.}}$ 48.84 67.86 23.21
$\bar{\phi}_{r}$ 58.91 82.14 30.36
TABLE V: Ablation study on DBSCAN and Canonical Query on scene "figurines" in LERF.
@@
SAM2 GT-anchored DBSCAN Cano. Query mIoU mAcc Loc. Acc
@@
V-C Ablation study.
We conduct ablation studies to validate the efficacy of each proposed component of our method and report the performance in Tab. IV on LERF’s "figurines" scene. Recall, $\bar{\phi}_{r}$ denotes the weighted average features. As a comparison, we also tested on a standard average strategy that is the sum of features from various views divided by the number of features, denoted as $\bar{\phi}_{r}^{\texttt{stand.}}$. As shown, mIoU increases significantly, i.e. +49.89, when using the GT-anchored query. In the meantime, SAM2’s masklets to get consistent features of the region also play an important role for the GT-anchored query to work.
Overall, we observe that all proposed components are necessary to achieve the optimal performance. We also studied the effectiveness of our method without DBSCAN [5] and evaluated the performance of canonical querying from LERF [12] on the task of 3D querying. The original definition of canonical query from LERF is conducted on rendered 2D images.
To evaluate the performance of canonical query in 3D querying, we replace $\phi_{\texttt{lang}}$ by ${\bm{l}}_{i}$ in the original equation. Formally, for each language embedding ${\bm{l}}_{i}$ and each text query $\bm{q}$}, the relevancy score is defined as $\min_{i}\frac{\exp({\bm{l}}_{i}\cdot{\bm{q}})}{\exp({\bm{l}}_{i}\cdot{\bm{q}})+\exp({\bm{q}}\cdot\phi_{canon}^{i})}$ where $\phi_{canon}^{i}$ is the CLIP embedding of a predefined canonical phrase chosen from “object", “things", “stuff", and “texture". The results are shown in Tab. V on LERF’s “figurines” scene. We see that without DBSCAN, our proposed method also gains significant improvement in performance. We also see that our proposed GT-anchored query significantly outperforms the canonical query.
VI Conclusion
@@
We study the task of open-vocabulary 3D understanding formulated as a point-level 3D querying task. Based on LangSplat’s framework, we present a tracking-based approach to provide consistent ground-truth supervision when distilling the language features. Furthermore, we introduce a novel GT-Anchored querying pipeline to address the difficulty of choosing a consistent threshold in the baseline’s single-step query. Experiments over three datasets demonstrate that our approach achieves state-of-the-art performance, with ablation studies verifying the efficacy of the proposed components.